OpenStudy (anonymous):

Let f(x)= {x+2 if x<3 {4x-7 if x>3 Which of the following statements are true about f? I. lim f(x) as x approaches 3 exists II. f is continuous at x=3 III. f is differentiable at x=3 the answer is I and II only, but how would you go about solving it to get the answer?

jigglypuff314 (jigglypuff314):

you could just plug in 3 for both equations and see that they intersect at x = 3 to do it limit wise... solve for the limits of \[\lim_{x \rightarrow 3^{-}} (x + 2)\] and \[\lim_{x \rightarrow 3^{+}} (4x - 7)\] 3^- part means as the limit as x is approaching 3 from the left which was taken from the x < 3 part

OpenStudy (anonymous):

x= 3 thn x+2 = 5 4x-7= 12-7=5 so f is continues as both connects in x=3 limit f(x) exists as both coming from negative and positive sides of 3 equals to 5 its not differentiable as there are two equations and takes a sharp turn at x=3

OpenStudy (anonymous):

@harindu how do you know it takes a sharp turn at x=3 without using a calculator?

OpenStudy (08surya):

for continuity left hand limit=right hand limit=value of the function at that point we will get left hand limit from x<3........so LHL=3+2=5 we will get right hand limit from x>3.....so RHL=4*3-7=5 and we will get value of function by puting x=3 in any expression value of function=3+2=5 so limit exist at x=3 bcause RHL=LHL and function is continuous becoz LHL=RHL=value of the function=3

OpenStudy (anonymous):

there are two functions and they connect at x=3 so it wont be a curve or a straight line so it would have a sharp turn